# Smooth polynomials with several prescribed coefficients

### v2

@@ 5.1. Major arc estimates
Let $40<\varepsilon<1/$ 4. For $a,b\in\mathbb{F}_{q}[t]$, $\gcd(a,g)=1$, $\ell+\deg g<n$, we have
@@
Let $40<\varepsilon<1/$ 4. For $b\in\mathbb{F}_{q}[t]$, $\gcd(b,t)=1$, $1\leq k<n-\ell$, we have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{bf}{t^{k}}\right)$ $=-\frac{1}{m}\frac{q^{n-\ell-k+1}}{q-1}\rho^{\prime}\left(\frac{n-k+1}{m}\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)$
$\quad+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{k}{2}}e^{\varepsilon k}\right),$
@@
$=\frac{1}{q^{\ell-1}(q-1)}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-k+1,m)$
$=\frac{q^{n-\ell-k+1}}{q-1}\left(\rho\left(\frac{n-k}{m}\right)-\rho\left(\frac{n-k+1}{m}\right)\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)$
by (3).
We have
$\rho\left(\frac{n-k}{m}\right)-\rho\left(\frac{n-k+1}{m}\right)$ $=-\frac{1}{m}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+\frac{\rho^{(2)}(\xi)}{2!}\frac{1}{m^{2}}$
for some $\xi\in((n-k)/m,(n-k+1)/m)$, thus by Lemmas 12 and 13 we have
$\rho\left(\frac{n-k}{m}\right)-\rho\left(\frac{n-k+1}{m}\right)$ $=-\frac{1}{m}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+O\left(\frac{1}{m^{2}}\rho^{(2)}\left(\frac{n-k+1}{m}\right)\right)$
$=-\frac{1}{m}\rho^{\prime}\left(\frac{n-k+1}{m}\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right).$
@@ 5.2. Minor arc estimates
#$\#\left\{f\in{\mathcal{M}}(n):\|\alpha f-\beta\|<q^{-k}\right\}\leq\max\left\{1,q^{n-k},q^{n}\|\alpha\|,q^{-k}/\|\alpha\|\right\}.$
@@
#$\#\left\{|e|=\frac{q^{i}}{|g|}:\|eg\gamma+\beta\|<q^{-j}\right\}\ll\max\left\{1,\frac{q^{i-j}}{|g|},\|g\gamma\|\frac{q^{i}}{|g|},\frac{1}{\|g\gamma\|q^{j}}\right\}$
@@ 6. Analysis of $S_{\mathcal{J}}$
Let $a,g\in\mathbb{F}_{q}[t]$ be two polynomials with $\gcd(a,g)=1$ and write $g=g_{0}t^{k}$, $\gcd(g_{0},t)=1$. If $11<|g_{0}|\leq q^{\lceil n/(\#{\mathcal{I}}+1)\rceil}-$ 1, then
@@ 7. Proof of Theorem 3
First we investigate the integral (9) over the major arcs defined by (10). We also choose
(27) $\kappa=\left\lceil\frac{n-m-2}{4}\right\rceil\quad\text{and}\quad\ell=\left\lceil\frac{n-m}{4}\right\rceil.$
@@
$M_{1}=\frac{1}{q^{n-\ell}}\sum_{0\leq k\leq\kappa}\sum_{\begin{subarray}{c}|a|<t^{k}\\ t\nmid a\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{t^{k}}\right)\right)$
@@
Concerning $M_{1}$, for $k=0$ we have by Lemma 18,
(29) $\begin{split}&\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\right)\\ =&\frac{1}{q^{n-\ell}}\left(\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right)\right)\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right).\end{split}$
Now consider the terms with $g=t^{k}$, $k\geq 1$. Let $B={\mathcal{I}}\cap\{n-\ell,\dots,n-1\}$. Clearly, one can represent the residue classes of ${\mathcal{R}}_{\ell}$ by the elements of ${\mathcal{M}}(\ell)$. The contribution of those $b$ that $b_{i-(n-\ell)}\neq\alpha_{i}$ for $i\in B$ is zero. Write
=${\mathcal{B}}=\left\{b\in{\mathcal{M}}(\ell):b_{j-(n-\ell)}=\alpha_{j}\text{ for }j\in B\right\}\quad\text{and}\quad I_{\ell}=I\cup\{n-\ell,\dots,n-1\}$
@@
$\beta_{i}=\left\{\begin{array}[]{cl}\alpha_{i}&\text{if $i\in{\mathcal{I}}$},\\ b_{i-(n-\ell)}&\text{if $i\in\{n-\ell,\dots,n-1\}\setminus I$ }.\end{array}\right.$
@@
#$\#{\mathcal{J}}_{b}=q^{n-\#{\mathcal{I}}_{\ell}}=q^{n-\#{\mathcal{I}}-\ell+\#B}$
@@
For $1\leq k\leq\kappa$ we recall that by Lemmas 12, 13 and 20, we have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)=-\frac{1}{m}\frac{q^{n-\ell-k+1}}{q-1}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+E(n,m,k)$
@@
$E(n,m,k)\ll q^{n-\ell-k}\rho\left(u\right)\frac{u^{1/2}(\log(u+1))^{5/2}}{m^{2}}+q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{k}{2}}e^{\varepsilon k}.$
By Lemma 24, $S_{{\mathcal{J}}_{b}}(a/t^{k})=0$ if $a_{i}\neq 0$ for some $k-i-1\not\in{\mathcal{I}}_{\ell}$, $0\leq i<k$. As $\gcd(a,t)=1$, we have $a_{0}\neq 0$ and thus it is enough to consider $k$ with $k-1\in{\mathcal{I}}_{\ell}$. For such a $k$, we have by Lemma 24 that
(30) $\begin{split}&\frac{1}{q^{n-\ell}}\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)S_{{\mathcal{J}}_{b}}(a/t^{k})\\ =&\frac{1}{q^{\#{\mathcal{I}}_{\ell}-\ell}}\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}_{\ell}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\sum_{\begin{subarray}{c}f\in S(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}\frac{a_{i}t^{i}f}{t^{k}}\right)\prod_{j\in{\mathcal{I}}_{\ell}}\mathbf{e}\left(\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}_{\ell}-\ell}}\left(-\frac{1}{m}\frac{q^{n-\ell-k+1}}{q-1}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+E(n,m,k)\right)\cdot S_{I}(n,k,\ell)\end{split}$
@@
$S_{I}(n,k,\ell)=\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}_{\ell}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\prod_{j\in{\mathcal{I}}_{\ell}}\mathbf{e}\left(\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right).$
Considering $S_{I}(n,k,\ell)$, as $k-i-1\leq k\leq\kappa<n-\ell$, for the summation over $a_{i}$ it is enough to consider $i$ with $k-i-1\in{\mathcal{I}}$. Moreover, for $j\geq n-\ell$ and $0\leq i<k$, we have $j+i-k\geq n-\ell-k>0$, and thus
@@
i.e., the terms in $S_{I}(n,k,\ell)$ do not depend on the choice of $b\in{\mathcal{B}}$. Therefore
$S_{I}(n,k,\ell)$ $=q^{\ell-\#B}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\prod_{j\in{\mathcal{I}}}\mathbf{e}\left(\alpha_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}\end{subarray}}a_{i}t^{i-k}\right)\right)$
@@
Write $I=\{i_{0}<i_{1}<\dots<i_{\nu}<\dots\}$ with $a_{i_{0}}=a_{i_{1}}=\dots=a_{i_{\nu-1}}=0\neq a_{i_{\nu}}$. Then taking summation over $0\leq k\leq\kappa$ we get by (29) and (30), that
(31) $\begin{split}M_{1}=&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right)\\ &+(q-1)\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<\kappa\end{subarray}}\frac{1}{q^{\#{\mathcal{I}}-j-\ell}}\left(-\frac{1}{m}\frac{q^{n-i_{j}-\ell}}{q-1}\rho^{\prime}\left(\frac{n-i_{j}}{m}\right)+E(n,m,i_{j}+1)\right)\\ &-\mathbf{1}(\nu<\kappa)\frac{1}{q^{\#{\mathcal{I}}-\nu-\ell}}\left(-\frac{1}{m}\frac{q^{n-i_{\nu}-\ell}}{q-1}\rho^{\prime}\left(\frac{n-i_{\nu}-\ell}{m}\right)+E(n,m,i_{\nu}+1)\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)\right)\\ &+O\left(\nu q^{n-\#{\mathcal{I}}-i_{0}-1}\rho\left(u\right)\frac{u^{1/2}(\log(u+1))^{5/2}}{m^{2}}+q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\frac{\kappa}{2}+\ell}e^{\varepsilon(n-m)/2}\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)\right)\\ &+O\left(q^{n-\#{\mathcal{I}}-i_{0}-1}\rho\left(u\right)\frac{u^{1/2}(\log(u+1))^{5/2}}{m}+q^{\left(\frac{7}{8}+\varepsilon\right)n-\frac{3m}{8}+\frac{5}{4}-\#{\mathcal{I}}}e^{\varepsilon(n-m)/2}\right)\end{split}$
by (27).
@@
$M_{2}=\frac{1}{q^{n-\ell}}\sum_{r=K}^{\kappa}\sum_{k=0}^{\kappa-r}M_{2}(r,k)$
@@
$\begin{split}M_{2}(r,k)&=\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{d\mid g_{0}t^{k}}\frac{\mu(g_{0}t^{k}/d)}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi_{g_{0}t^{k}/d}(n-\deg d,m)+R(r,k)\right)\\ &\qquad\cdot\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)\end{split}$
@@
By (12) and Lemmas 12 and 15, we have
@@
$\ll$ $\sum_{d\mid g_{0}t^{k}}\frac{q^{n-\deg d}}{q^{\ell}\Phi(g_{0}t^{k}/d)}\rho\left(\frac{n-\deg d}{m}\right)$
$\ll$ $q^{n-\ell}\rho\left(u\right)\sum_{d\mid g_{0}t^{k}}\frac{(u\log u)^{\deg d/m}}{|d|\Phi(g_{0}t^{k}/d)}$
$\ll$ $q^{n-\ell-r-k}\rho\left(u\right)\left(u\log u\right)^{1/8}(r+k)\left(1+\log_{q}(r+k)\right).$

### v3

@@ 5.1. Major arc estimates
Let $0<\varepsilon<1/4$. For $a,b\in\mathbb{F}_{q}[t]$, $\gcd(a,g)=1$, $\ell+\deg g<n$, we have
@@
Let $0<\varepsilon<1/4$. For $b\in\mathbb{F}_{q}[t]$, $\gcd(b,t)=1$, $1\leq k<n$-$\ell$, we have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{bf}{t^{k}}\right)=$ $\frac{1}{q^{\ell-1}(q-1)}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-k+1,m)$
$+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n+k/2}e^{\varepsilon(\ell+k}\right),$
@@
$=\frac{1}{q^{\ell-1}(q-1)}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-k+1,m).$
@@ 5.2. Minor arc estimates
$\left\{f\in{\mathcal{M}}(n):\|\alpha f-\beta\|<q^{-k}\right\}\leq\max\left\{1,q^{n-k},q^{n}\|\alpha\|,q^{-k}/\|\alpha\|\right\}.$
@@
$\left\{|e|=\frac{q^{i}}{|g|}:\|eg\gamma+\beta\|<q^{-j}\right\}\ll\max\left\{1,\frac{q^{i-j}}{|g|},\|g\gamma\|\frac{q^{i}}{|g|},\frac{1}{\|g\gamma\|q^{j}}\right\}$
@@ 6. Analysis of $S_{\mathcal{J}}$
Let $a,g\in\mathbb{F}_{q}[t]$ be two polynomials with $\gcd(a,g)=1$ and write $g=g_{0}t^{k}$, $\gcd(g_{0},t)=1$. If $1<|g_{0}|\leq q^{\lceil n/(\#{\mathcal{I}}+1)\rceil}-1$, then
@@ 7. Proof of Theorem 4
First, we investigate the integral (17) over the major arcs defined by (18).
@@
$M_{1}=\sum_{0\leq k\leq\kappa}M_{1,k}$
where
$M_{1,k}=\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}|a|<t^{k}\\ t\nmid a\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{t^{k}}\right)\right)$
@@
Concerning $M_{1,0}$, we have by Lemma 19,
(35) $\begin{split}M_{1,0}=&\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\right)\\ =&\frac{1}{q^{n-\ell}}\left(\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right)\right)\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right).\end{split}$
Now consider the terms with $M_{1,k}$ with $k\geq 1$. Let $B={\mathcal{I}}\cap\{n-\ell,\dots,n-1\}$. Clearly, one can represent the residue classes of $\mathcal{R}$$_{\ell}$ by the elements of ${\mathcal{M}}(\ell)$. The contribution of those $b$ that $b_{i-(n-\ell)}\neq\alpha_{i}$ for $i\in B$ is zero. Write
${\mathcal{B}}=\left\{b\in{\mathcal{M}}(\ell):b_{j-(n-\ell)}=\alpha_{j}\text{ for }j\in B\right\}\quad\text{and}\quad{\mathcal{I}}_{\ell}={\mathcal{I}}\cup\{n-\ell,\dots,n-1\}$
@@
$\beta_{i}=\left\{\begin{array}[]{cl}\alpha_{i}&\text{if $i\in{\mathcal{I}}$},\\ b_{i-(n-\ell)}&\text{if $i\in\{n-\ell,\dots,n-1\}\setminus{\mathcal{I}}$ }.\end{array}\right$.
@@
(36) ${\mathcal{J}}_{b}=q^{n-\#{\mathcal{I}}_{\ell}}=q^{n-\#{\mathcal{I}}-\ell+\#B}$
@@
Whence
$M_{1,k}=\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}|a|<t^{k}\\ t\nmid a\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)\overline{S_{{\mathcal{J}}_{b}}\left(a/t^{k}\right)}\right)$
For $1\leq k\leq\kappa$, we recall that by Lemma 21, we have
(37) $\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)=\frac{1}{q^{\ell-1}(q-1)}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-k+1,m)+E(n,m,k),$
@@
$E(n,m,k)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n+k/2}e^{\varepsilon(\ell+k)}.$
By Lemma 25, $S_{{\mathcal{J}}_{b}}(a/t^{k})=0$ if $a_{i}\neq 0$ for some $k-i-1\not\in{\mathcal{I}}_{\ell}$, $0\leq i<k$. As $\gcd(a,t)=1$, we have $a_{0}\neq 0$ and thus it is enough to consider $k$ with
(38) $k-1\in{\mathcal{I}}_{\ell}.$
For such a $k$, we have by Lemma 25 and by (37) that
(39) $\begin{split}M_{1,k}=&\frac{1}{q^{n-\ell}}\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)\overline{S_{{\mathcal{J}}_{b}}(a/t^{k})}\\ =&\frac{1}{q^{\#{\mathcal{I}}_{\ell}-\ell}}\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}_{\ell}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\sum_{\begin{subarray}{c}f\in S(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}\frac{a_{i}t^{i}f}{t^{k}}\right)\prod_{j\in{\mathcal{I}}_{\ell}}\mathbf{e}\left(-\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}_{\ell}-\ell}}\left(\frac{1}{q^{\ell-1}(q-1)}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-k+1,m)+E(n,m,k)\right)\cdot\overline{S_{\mathcal{I}}(n,k,\ell)}\end{split}$
@@
$S_{\mathcal{I}}(n,k,\ell)=\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}_{\ell}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\prod_{j\in{\mathcal{I}}_{\ell}}\mathbf{e}\left(\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right).$
Considering $S_{\mathcal{I}}(n,k,\ell)$, as $k-i-1\leq k\leq\kappa<n$-$\ell$, for the summation over $a_{i}$ it is enough to consider $i$ with $k-i-1\in{\mathcal{I}}$. Moreover, for $j\geq n$-$\ell$ and $0\leq i<k$, we have $j+i-k\geq n-\ell-k>0$, and thus
@@
i.e., the terms in $S_{\mathcal{I}}(n,k,\ell)$ do not depend on the choice of $b\in{\mathcal{B}}$. Therefore
(40) $S_{\mathcal{I}}(n,k,\ell)$ $=q^{\ell-\#B}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\prod_{j\in{\mathcal{I}}}\mathbf{e}\left(\alpha_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}\end{subarray}}a_{i}t^{i-k}\right)\right)$
@@
Write $I=\{i_{0}<i_{1}<\dots<i_{\nu}<\dots\}$ with $a_{i_{0}}=a_{i_{1}}=\dots=a_{i_{\nu-1}}=0\neq a_{i_{\nu}}$. Then taking summation over $0\leq k\leq\kappa$ we get by (35), (36), (38), (39) and (40), that
(41) $\begin{split}M_{1}=&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right)+\sum_{\begin{subarray}{c}1\leq k\leq\kappa\\ k-1\in{\mathcal{I}}_{\ell}\end{subarray}}M_{1,k}\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right)\\ &+(q-1)\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<\kappa\end{subarray}}\frac{1}{q^{\#{\mathcal{I}}-j-\ell}}\Bigg{(}\frac{1}{q^{\ell-1}(q-1)}\Psi(n-i_{j}-1,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-i_{j},m)\\ &\qquad\qquad\qquad\qquad\qquad\qquad+E(n,m,i_{j}+1)\Bigg{)}\\ &-\mathbf{1}(i_{\nu}<\kappa)\frac{1}{q^{\#{\mathcal{I}}-\nu-\ell}}\Bigg{(}\frac{1}{q^{\ell-1}(q-1)}\Psi(n-i_{\nu}-1,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-i_{\nu},m)\\ &\qquad\qquad\qquad\qquad\qquad\qquad+E(n,m,i_{\nu}+1)\Bigg{)}\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}})+O\left(q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa/2+\ell+\nu}e^{\varepsilon(\ell+\kappa)}\right)\right).\end{split}$
@@
(42) $M_{2}=\frac{1}{q^{n-\ell}}\sum_{r=K}^{\kappa}\sum_{k=0}^{\kappa-r}M_{2}(r,k)$
@@
$\begin{split}M_{2}(r,k)&=\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{d\mid g_{0}t^{k}}\frac{\mu(g_{0}t^{k}/d)}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi_{g_{0}t^{k}/d}(n-\deg d,m)+R(r,k)\right)\\ &\qquad\cdot\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)\\ &\ll\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\left(\sum_{d\mid g_{0}t^{k}}\frac{|\mu(g_{0}t^{k}/d)|}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi_{g_{0}t^{k}/d}(n-\deg d,m)+R(r,k)\right)\\ &\qquad\cdot\left|\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}S_{J}\left(\frac{a}{g_{0}t^{k}}\right)\right|\end{split}$
@@
By (4), (20) and Lemma 14, we have
$\sum_{d\mid g_{0}t^{k}}\frac{|\mu(g_{0}t^{k}/d)|}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi_{g_{0}t^{k}/d}(n-\deg d,m)$
$\ll$ $\sum_{d\mid g_{0}t^{k}}\frac{q^{n-\deg d}}{q^{\ell}\Phi(g_{0}t^{k}/d)}\rho\left(\frac{n-\deg d}{m}-\frac{n-\deg d}{2m^{2}}\right)\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)$
$\ll$ $q^{n-\ell}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\sum_{d\mid g_{0}t^{k}}\frac{(u\log u)^{\deg d/m}}{|d|\Phi(g_{0}t^{k}/d)}$
$\ll$ $q^{n-\ell-r-k}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)$
$\qquad\cdot\left(u\log u\right)^{(r+k)/m}\left(1+\log_{q}(r+k)\right)\tau(g_{0}t^{k}).$
Lemma  16 yields $\tau(g_{0}t^{k})\ll q^{(2+\varepsilon)(r+k)/\log(r+k)}\ll q^{(r+k)/5}$. Moreover,
$\left(u\log u\right)^{(r+k)/m}\leq\exp\left\{\frac{11\log u}{10}\cdot\frac{r+k}{m}\right\}\leq\exp\left\{\frac{11(r+k)}{10}\cdot\frac{\log n}{m}\right\}\leq q^{3\frac{r+k}{5}}$
by (6) and (7). Thus
@@
$\ll q^{n-\ell-(r+k)/5}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\left(1+\log_{q}(r+k)\right)$
with error term
$T(n,m)\ll_{\varepsilon}\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}$